11% □% (Round to two decimal places.)
Answer #1

Return of portfolio = 27% * 13.33% + 47% * 12.33% + 26% * 6.66%

                                   = 3.59% + 5.79% + 1.73%

                                   = 11.11%

Working note:

Return of portfolio in boom = (1 / 3) * 13% + (1 / 3) * 22% + (1 / 3) * 5%

                                              = 13.33%

Return of portfolio in normal = (1 / 3) * 11% + (1 / 3) * 15% + (1 / 3) * 11%

                                                = 12.33%

Return of portfolio in recession = (1 / 3) * 5% + (1 / 3) * 2% + (1 / 3) * 13%

                                                    = 6.66%